OpenStudy (anonymous):

The limit limh→0(6+h)2−36/h represents the derivative of some function f(x) at some number a. Find f and a.

OpenStudy (anonymous):

((f(x+h)-f(x))/h f(x)=x^2 x=6 so (((x+6)^2-36)/h it's drivation formules so f(x)'=2x f(6)'=12

OpenStudy (anonymous):

and what A will be

OpenStudy (anonymous):

i think a is 6 x=a=6

OpenStudy (anonymous):

thx